OpenStudy (anonymous):

Simplify: 4 x 3^2 - 5 + 2 x 3 - 11

OpenStudy (skullpatrol):

any ideas?

OpenStudy (skullpatrol):

have you studied the order of operations?

OpenStudy (skullpatrol):

4*3^2 - 5 + 2*3 - 11

OpenStudy (skullpatrol):

1. first simplify 3^2

OpenStudy (skullpatrol):

3^2 = 3*3

OpenStudy (anonymous):

9

OpenStudy (anonymous):

PEMDAS kinda confuses me

OpenStudy (anonymous):

Do i multiply first?

OpenStudy (skullpatrol):

replace 3^2 with 9

OpenStudy (anonymous):

ok

OpenStudy (skullpatrol):

4*3^2 - 5 + 2*3 - 11 4*9 - 5 + 2*3 - 11

OpenStudy (anonymous):

Is it 36-5+6-11 now?

OpenStudy (skullpatrol):

Do all the additions and subtractions in order from left to right.

OpenStudy (anonymous):

I got 36

OpenStudy (skullpatrol):

How?

OpenStudy (anonymous):

36-5+6-11

OpenStudy (skullpatrol):

36 - 5 = ?

OpenStudy (anonymous):

31

OpenStudy (skullpatrol):

replace 36 - 5 with 31

OpenStudy (anonymous):

oh wait, never mind. I subtracted 1 not 11.

OpenStudy (anonymous):

its 26

OpenStudy (anonymous):

thanks

OpenStudy (skullpatrol):

thanks for asking :-)